# Nontriviality of rings of integral-valued polynomials

### v1

@@ 5 Ramification Indices and Residue Field Degrees
$\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})\text{ is nontrivial }\stackrel{{\scriptstyle\text{Prop}3.9}}{{\Longleftrightarrow}}\textnormal{Int}_{\mathbb{Q}}(\mathcal{S}_{p},\overline{\mathbb{Z}_{p}})\text{is nontrivial }\stackrel{{\scriptstyle\text{Prop}3.10}}{{\Longleftrightarrow}}\textnormal{Int}_{\overline{\mathbb{Q}_{p}}}(\mathcal{S}_{p},\overline{\mathbb{Z}_{p}})\text{ is nontrivial }$
@@
    {#$\{\#S/\overline{P}\mid\overline{P}\in\mathcal{P}_{p}\}$ is unbounded.
@@
[…]
@@
Moreover, for each $\overline{P}\in\mathcal{P}_{p}$, $\tau_{\overline{P}}$ establishes a bijection between $S/\overline{P}$ and $\tau_{\overline{P}}(S)/\overline{M_{p}}$. If $\tau_{\overline{P}}(S)/\overline{M_{p}}$ is infinite for some $\overline{P}\in\mathcal{P}_{p}$, then considering the pullback via $\tau_{\overline{P}}^{-1}$ we have that $S/\overline{P}$ is infinite. If instead $\tau_{\overline{P}}(S)/\overline{M_{p}}$ is finite for each $\overline{P}\in\mathcal{P}_{p}$, since $\mathcal{S}_{p}/\overline{M_{p}}$ is infinite, this means that there exists a sequence $\{\overline{P}_{n}\}_{n\in\mathbb{N}}\subseteq\mathcal{P}_{p}$ such that #$\#\tau_{\overline{P}_{n}}(S)/\overline{M_{p}}=d_{n}<\infty$ is diverging to infinity. This amounts to say via the previous one-to-one correspondence that #$\#S/\overline{P}_{n}=d_{n}$ is diverging to infinity, which means that {#$\{\#S/\overline{P}\mid\overline{P}\in\mathcal{P}_{p}\}$ is unbounded.
@@
For this domain $D$, every valuation overring is a DVR, and every residue field $D$ is equal to $\mathbb{F}_{p}$. These conditions guarantee that the ramification indices of $D$ are bounded. Furthermore, $D$ cannot contain any pseudo-divergent sequence with respect to any of its maximal prime ideals $\overline{P}\cap D$, because $D_{\overline{P}\cap D}$ is a DVR. We will show that $\mathcal{D}_{p}^{\prime}:=\bigcup_{\overline{P}\in\mathcal{P}_{p}}\tau_{\overline{P}}(D_{\overline{P}\cap D})\subseteq\overline{\mathbb{Z}_{p}}$ contains a pseudo-divergent sequence and by an argument similar to Lemma 5.11, the ring $\textnormal{Int}_{\mathbb{Q}}(D,\overline{\mathbb{Z}_{(p)}})$ must equal $\mathbb{Z}_{(p)}[X]$.
@@
    For each $n\in\mathbb{N}$, there exist $P_{n}\subset D$, $p\in P_{n}$ and $s_{n}\in D_{P_{n}}$ such that $0<v_{P_{n}}(s_{n})<\frac{1}{n}$.
@@
(3) $\Rightarrow$ (1) Assume that (3) holds. Then, there exists a pseudo-divergent sequence $E=\{\tilde{s}_{n}\}_{n\in\mathbb{N}}\subset\mathcal{D}_{p}^{\prime}$ with $\textnormal{Br}(E)=\overline{M_{p}}$. Hence, $v_{p}(\tilde{s}_{n})$ strictly decreases to 00 and, up to taking a subsequence, we may assume that $0<v_{p}(\tilde{s}_{n})<\frac{1}{n}$. Since $\tilde{s}_{n}=\tau_{P_{n}}(s_{n})$ for some $P_{n}\subset D$, $p\in D$ and $s_{n}\in D_{P_{n}}$, we have $0<v_{P_{n}}(s_{n})=v_{p}(\tau_{P_{n}}(s_{n}))<\frac{1}{n}$ which implies that $n<e(P_{n}\mid p)$. Thus $\{e(P_{n}\mid p)\}_{n\in\mathbb{N}}$ is unbounded. Note that, as in the proof of (2) $\Rightarrow$ (3), if $P_{n}$ is the same prime ideal $P$ for infinitely many $n$’s, then $e(P\mid p)=\infty$.
@@
(3) $\Rightarrow$ (1) Let $K^{\prime}$ be the Galois closure of $K$ and let $D^{\prime}$ be the integral closure of $D$ in $K^{\prime}$. By Lemma 5.17, $E_{D^{\prime},p}$ and $F_{D^{\prime},p}$ are bounded. Therefore, $\textnormal{Int}_{K^{\prime}}(D^{\prime})$ is Prüfer by [14, Theorem 2.5]. We consider now the set $G(D)=\bigcup_{\sigma\in G}\sigma(D)\subset D^{\prime}$. Since $\textnormal{Int}_{K^{\prime}}(D^{\prime})\subseteq\textnormal{Int}_{K^{\prime}}(G(D),D^{\prime})$, the latter ring is Prüfer, too. By [22, Lemma 2.20], $\textnormal{Int}_{K^{\prime}}(G(D),D^{\prime})$ is the integral closure in $K^{\prime}(X)$ of $\textnormal{Int}_{\mathbb{Q}}(D)=\textnormal{Int}_{\mathbb{Q}}(G(D),D^{\prime})$. It follows by [11, Theorem 22.4] that $\textnormal{Int}_{\mathbb{Q}}(D)=\textnormal{Int}_{K^{\prime}}(G(D),D^{\prime})\cap\mathbb{Q}(X)$ is Prüfer. ∎
@@ 6 Polynomial closure of subsets of algebraic integers
Let $S\subseteq\overline{\mathbb{Z}}$. Let $\{f_{n}(X)/d_{n}\}_{n\geq 0}$ be a regular basis for $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ such that each $f_{n}\in\mathbb{Z}[X]$ is monic and each $d_{n}$ is a positive integer. Let =$\mathscr{F}=\{f_{n}(X)/d_{n}\mid d_{n}\geq 2\}$.

### v2

@@ 5 Ramification Indices and Residue Field Degrees
[…]
@@
    $\{\#S/\overline{P}\mid\overline{P}\in\mathcal{P}_{p}\}$ is unbounded.
@@
(1) $\Rightarrow$ (2). Suppose that $\{\#S/\overline{P}\mid\overline{P}\in\mathcal{P}_{p}\}$ is unbounded, but $\mathcal{S}_{p}/\overline{M_{p}}=n$ for some $n\in\mathbb{N}$. By assumption, there exists some $\overline{P}\in\mathcal{P}_{p}$ such that $S/\overline{P}>n$. Thus, there exist $s_{0},\ldots,s_{n}\in S$ such that $u_{\overline{P}}(s_{i}-s_{j})=0$ for each $0\leq i<j\leq n$. Applying the $\mathbb{Q}$-embedding $\tau_{\overline{P}}$ to each $s_{i}$, we obtain $n+1$ elements $\tau_{\overline{P}}(s_{0}),\ldots,\tau_{\overline{P}}(s_{n})\in\mathcal{S}_{p}$ such that
@@
Moreover, for each $\overline{P}\in\mathcal{P}_{p}$, $\tau_{\overline{P}}$ establishes a bijection between $S/\overline{P}$ and $\tau_{\overline{P}}(S)/\overline{M_{p}}$. If $\tau_{\overline{P}}(S)/\overline{M_{p}}$ is infinite for some $\overline{P}\in\mathcal{P}_{p}$, then considering the pullback via $\tau_{\overline{P}}^{-1}$ we have that $S/\overline{P}$ is infinite. If instead $\tau_{\overline{P}}(S)/\overline{M_{p}}$ is finite for each $\overline{P}\in\mathcal{P}_{p}$, since $\mathcal{S}_{p}/\overline{M_{p}}$ is infinite, this means that there exists a sequence $\{\overline{P}_{n}\}_{n\in\mathbb{N}}\subseteq\mathcal{P}_{p}$ such that $\tau_{\overline{P}_{n}}(S)/\overline{M_{p}}=d_{n}<\infty$ is diverging to infinity. This amounts to say via the previous one-to-one correspondence that $S/\overline{P}_{n}=d_{n}$ is diverging to infinity, which means that $\{\#S/\overline{P}\mid\overline{P}\in\mathcal{P}_{p}\}$ is unbounded.
@@
For this domain $D$, every valuation overring is a DVR, and every residue field of $D$ is equal to $\mathbb{F}_{p}$. These conditions guarantee that the ramification indices of $D$ are bounded. Furthermore, $D$ cannot contain any pseudo-divergent sequence with respect to any of its maximal prime ideals $\overline{P}\cap D$, because $D_{\overline{P}\cap D}$ is a DVR. We will show that $\mathcal{D}_{p}^{\prime}:=\bigcup_{\overline{P}\in\mathcal{P}_{p}}\tau_{\overline{P}}(D_{\overline{P}\cap D})\subseteq\overline{\mathbb{Z}_{p}}$ contains a pseudo-divergent sequence and by an argument similar to Lemma 5.11, the ring $\textnormal{Int}_{\mathbb{Q}}(D,\overline{\mathbb{Z}_{(p)}})$ must equal $\mathbb{Z}_{(p)}[X]$.
@@
    For each $n\in\mathbb{N}$, there exist $P_{n}\subset D$, with $p\in P_{n}$ and $s_{n}\in D_{P_{n}}$ such that $0<v_{P_{n}}(s_{n})<\frac{1}{n}$.
@@
(3) $\Rightarrow$ (1) Assume that (3) holds. Then, there exists a pseudo-divergent sequence $E=\{\tilde{s}_{n}\}_{n\in\mathbb{N}}\subset\mathcal{D}_{p}^{\prime}$ with $\textnormal{Br}(E)=\overline{M_{p}}$. Hence, $v_{p}(\tilde{s}_{n})$ strictly decreases to 0 and, up to taking a subsequence, we may assume that $0<v_{p}(\tilde{s}_{n})<\frac{1}{n}$. Since $\tilde{s}_{n}=\tau_{P_{n}}(s_{n})$ for some $P_{n}\subset D$, $p\in D$ and $s_{n}\in D_{P_{n}}$, we have $0<v_{P_{n}}(s_{n})=v_{p}(\tau_{P_{n}}(s_{n}))<\frac{1}{n}$ which implies that $n<e(P_{n}\mid p)$. Thus $\{e(P_{n}\mid p)\}_{n\in\mathbb{N}}$ is unbounded. Note that, as in the proof of (2) $\Rightarrow$ (3), if $P_{n}$ is the same prime ideal $P$ for infinitely many $n$’s, then $e(P\mid p)=\infty$.
@@
(3) $\Rightarrow$ (1) Let $K^{\prime}$ be the Galois closure of $K$ and let $D^{\prime}$ be the integral closure of $D$ in $K^{\prime}$. By Lemma 5.17, $E_{D^{\prime},p}$ and $F_{D^{\prime},p}$ are bounded. Therefore, $\textnormal{Int}_{K^{\prime}}(D^{\prime})$ is Prüfer by [14, Theorem 2.5]. We consider now the set $G(D)=\bigcup_{\sigma\in G}\sigma(D)\subset D^{\prime}$. Since $\textnormal{Int}_{K^{\prime}}(D^{\prime})\subseteq\textnormal{Int}_{K^{\prime}}(G(D),D^{\prime})$, the latter ring is Prüfer, too. By [23, Lemma 2.20], $\textnormal{Int}_{K^{\prime}}(G(D),D^{\prime})$ is the integral closure in $K^{\prime}(X)$ of $\textnormal{Int}_{\mathbb{Q}}(D)=\textnormal{Int}_{\mathbb{Q}}(G(D),D^{\prime})$. It follows by [11, Theorem 22.4] that $\textnormal{Int}_{\mathbb{Q}}(D)=\textnormal{Int}_{K^{\prime}}(G(D),D^{\prime})\cap\mathbb{Q}(X)$ is Prüfer. ∎
@@ 6 Polynomial closure of subsets of algebraic integers
Let $S\subseteq\overline{\mathbb{Z}}$. Let $\{f_{n}(X)/d_{n}\}_{n\geq 0}$ be a regular basis for $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ such that each $f_{n}\in\mathbb{Z}[X]$ is monic and each $d_{n}$ is a positive integer. Let $\mathscr{F}=\{f_{n}(X)/d_{n}\mid d_{n}\geq 2\}$.